Circular No. 8384

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004dq T. Boles, Coddenham, England, reports the discovery of an apparent

supernova (mag 18.5) on unfiltered CCD images taken on June 27.929 and 28.942 UT with a 0.35-m reflector. The new object is located at α = 15^{h}44^{m}46^{s}.84, δ = +22^{o}50′22′′.9, which is ≈ 4′′.9 east and 4′′.8 north of the center of its apparent host galaxy; measurement by Boles of Sloan-r′ images taken by M. Bode, J. Marchant, and A. Newsam on Aug. 9.959 with the 2-m Liverpool telescope at La Palma yields position end figures 46^{s}.89, 23′′.0 (and mag ∼ 18.3) for SN 2004dq. The new object is not present on Boles’ images taken on 2004 Mar. 20 and May 18 (limiting mag 19.5) and on Digitized Sky Survey images (limiting red mag 22.0; limiting blue mag 21.5).

SUPERNOVA 2004dr IN ESO 479-G42 Further to IAUC 8381, J. Graham and W. Li report the LOSS discovery

of an apparent supernova (mag ∼ 17.5) on unfiltered KAIT images taken on Aug. 9.45 and 10.47 UT. The new object was also marginally detected on an earlier image taken on July 30.44 (limiting mag ∼ 18.5). SN 2004dr is located at α = 2^{h}46^{m}53^{s}.97, δ = −22^{o}38′15′′.0 (equinox 2000.0), which is 7′′.7 east and 31′′.6 north of the nucleus of ESO 479-G42. A KAIT image taken on Jan. 20.21 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2004dj IN NGC 2403 E. W. Guenther and S. Klose, Thüringer Landessternwarte, Tauten-

burg (TLT), report that observations of SN 2004dj, obtained with the Tautenburg 2-m telescope (+ high-resolution échelle spectrograph) on Aug. 4.08 and 5.06 UT, reveal Na I D lines in NGC 2403 with a heliocentric velocity of 164.8 ± 0.1 km/s (each with equivalent width 0.092 ± 0.005 Ångström). Following Munari and Zwitter (1997, A.Ap. 318, 269), this corresponds to E(B − V) = 0.026 ± 0.002 mag of line-of-sight reddening in the host.

K. Weis and D. J. Bomans, Astronomisches Institut, Ruhr-Universität, Bochum; and S. Klose, TLT, report identification of the potential progenitor of SN 2004dj (IAUC 8377, 8378, 8379) as an Hα-bright, pointlike source on CCD images taken with the Tautenburg 2-m telescope in April 2003. The position of this source is consistent with that of the luminous-blue-variable candidate 7-3909 identified by F. Spiller (1992, Ph.D. thesis, Heidelberg) at α = 7^{h}37^{m}16^{s}.93, δ = +65^{o}35′57′′.6 (equinox 2000.0), with V = 18.9, based on Calar Alto 3.5-m telescope data taken in February 1990.

2004 August 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green